After being drafted in the first round of the NFL draft, a star defensive end invests his signing bonus of $9,120,500.00 in a mutual fund. The fund pays on average 7.00% APR. The player will not touch this money until he retires from the league in 15.00 years. How much will this “nest egg” be worth at retirement?

A very careful new father wants to set money aside for his baby daughter’s wedding. If the wedding takes place in 23.00 years, he expects it will cost $88,500.00 . If the father can earn 5.00% on his investments, how much does he need to invest today to reach his goal?

An engaged couple agrees to get married once their savings equal $53,200.00. They have $32,300.00 at the moment and will invest the money in an account that earns 7.00% per year. How many years will the engagement last?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

1.

A=$9,120,500*(1.07)^15

=$9,120,500*2.759031541

=$25,163,747.17(Approx).

2.

88500=P*(1.05)^23

P=88500/1.05^23

=88500*0.325571305

=$28813.06(Approx).

3.

53200=32300*(1.07)^n

(53200/32300)=(1.07)^n

Taking log on both sides;

log (53200/32300)=n*log 1.07

n=log(53200/32300)/log 1.07

=7.38 years(Approx).